### v1

# Self interacting scalar field theory in general curved spacetimes at zero and finite temperature revisited

Abstract
We revisit the problem of spontaneous symmetry breaking (SSB), its restoration and phase transition for a self interacting quantum scalar field in a general curved background, at zero and finite temperature. To the best of our knowledge, most of the earlier computations in this context have been done in the linear order in curvature, which may not be very suitable for the Ricci flat spacetimes. One of our objectives is to see whether the higher order terms can bring in qualitatively new physical effects, and thereby attempting to fill in this gap in the literature. We use the Bunch-Parker local momentum space representation of the Feynman propagator. We compute the renormalised, background spacetime curvature (up to quadratic order) and temperature dependent one loop effective potential for $\phi^{4}$ plus $\phi^{3}$ self interaction. In particular for the de Sitter spacetime, we have shown for $\phi^{4}$-theory that we can have SSB for a scalar even with a positive rest mass squared and non-minimal coupling, at zero temperature. This cannot be achieved by the linear curvature term alone and the result remains valid for a very large range of renormalisation scale. For a phase transition, we have computed the leading curvature correction to the critical temperature. At finite temperature, symmetry restoration is also demonstrated. We also extend some of the above results to two loop level. The symmetry breaking in de Sitter at two loop remains present. We have further motivated the necessity of treating this problem non-perturbatively in some instances.
Keywords : Thermal field theory, curved spacetime, effective potential, symmetry breaking, de Sitter
1 Introduction
The motivation of studying quantum field theory in curved backgrounds is to investigate the effects of spacetime curvature or geometry on the fluctuations or propagation of quantum fields living in it. This much cultivated yet still active subject, originally supposed to be a prelude to a full theory of quantum gravity, has led to interesting physical predictions like the particle creation in cosmological or black hole backgrounds, fluctuations in the early inflationary universe, regularisation of the stress energy momentum tensor and so on, over decades. We refer our reader to [1, 2, 3, 4] and references therein for a vast historic overview.
We are interested to look into a self interacting scalar field theory at zero as well as at finite temperature in a general, regular curved background in this paper. Thermal fields in a curved background may exist in various different scenarios. For instance, the thermal fields living in the early universe, or the hot plasma residing in a star or accreting onto a black hole, the blackbody radiation emitted by a black hole, and so on.
An interesting aspect of quantum field theory with a background or classical scalar field is the existence of an effective potential. Such potential arises due to the loop integration over the quantum fluctuations and it adds up to the tree level self interaction. The total effective potential (i.e., the tree plus the loop correction) can be qualitatively different in shape from that of the tree level. Such effective potential may lead to the highly interesting physical phenomenon of spontaneous symmetry breaking, developed in the seminal works in [5, 6, 7, 8]. We further refer our reader to [9, 10, 11, 12] and references therein for various aspects associated with symmetry breaking and effective potentials. Computation of effective potentials for field theories at finite temperature is also a physically well motivated problem. Interestingly, the effective potential at high temperature ceases to have any symmetry breaking feature even if at zero temperature it has so, known as the symmetry restoration, e.g. [7, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. We further refer our reader to [25, 26, 27] and references therein for vast discussion on these issues.
We do not have the scope here to review in detail the developments of quantum field theory in curved spacetimes, and the effective action formalism. We refer our reader to [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59] and references therein for some aspects of propagators, renormalisation, effective action and symmetry breaking issues at zero temperature. For thermal propagators and finite temperature effects in curved spacetime, we refer our reader to [60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78] and also references therein. As far as symmetry breaking and phase transition in a general curved background is concerned, most of the earlier works use linear in the curvature approximation and then employ some renormalisation group technique for resummation (e.g. [33, 35, 38, 40] and also references therein). For example in [38], a renormalisation group derivation of the effective potential for a massless scalar coupled to gauge field, and discussion on phase transition can be seen. Even though this is a powerful method, it does not work for the Ricci flat spacetimes. In particular, the $R\phi^{2}$-type non-minimal coupling does not contribute to the effective potential for such spacetimes. At finite temperature on the other hand, the Page approximation technique can be useful for static spacetimes [62]. To the best of our knowledge, the thermal effective Lagrangian density in a general curved spacetime for $\phi^{4}$ theory, up to quadratic order in the curvature was first derived in [67], by employing a proper time expansion technique.
We wish to revisit in this paper the issue of the effective potential for a self interacting scalar field theory at both zero as well as finite temperature. We wish to put an emphasis on the spontaneous symmetry breaking and its restoration and phase transition phenomena. We shall keep curvature terms up to quadratic order containing the Kretschmann scalar and hence our result will also be valid for the Ricci flat spacetimes. One of our objectives behind this work is to see whether the higher order curvature terms could bring in qualitatively new effects, and thereby attempting to fill in the gap that seems to be present in the literature. We shall also discuss generalisation of some of the one loop results to two loop level.
The rest of the paper is organised as follows. In the next section, we discuss very briefly the basic technical framework we work in. In 3, we briefly derive the one loop effective potential at zero temperature, up to the quadratic order of the spacetime curvature for quartic and cubic self interactions. In 4, we extend these results to finite temperature, with high temperature approximation. Although as we mentioned, the above problem was first addressed in [67], our notation and the choice of the renormalisation scale will be a bit different. Due to this reason, we shall keep briefly the computations, for the sake of completeness and clarity. We discuss spontaneous symmetry breaking, its restoration and phase transition at zero and finite temperatures in 5, 5.1. We next generalise some of these results to two loop in 6. In particular, the contribution from the two loop double bubble diagram under local approximation to the effective potential, at finite temperature has been explicitly presented, 6.1. Finally we conclude in 7 with some future motivation for non-perturbative calculations. For the de Sitter spacetime in particular, we have shown that we can have symmetry breaking for a scalar even with positive rest mass squared and a positive non-minimal coupling, at zero temperature, at both one and two loop level. This cannot be achieved by the linear curvature term alone and the result remains valid for a very large range of renormalisation scale. Also, we have obtained this result with a cosmological constant as tiny as its current observed value.
We will work with the mostly positive signature of the metric in $d=4-\epsilon$ ($\epsilon=0^{+}$) and will set $c=\hbar=k_{B}=1$ throughout.
2 The basic set up
We wish to begin by briefly reviewing below the basic framework we will be working in, referring our reader to [1, 4, 32, 34] and references therein for detail. The action of the self interacting scalar field theory we are interested in reads in a curved spacetime background
$S=-\int d^{d}x\sqrt{-g}\left[\frac{1}{2}g^{\mu\nu}(\nabla_{\mu}\Phi)(\nabla_{\nu}\Phi)+\frac{1}{2}m^{2}\Phi^{2}+\frac{\eta\Phi^{3}}{3!}+\frac{\lambda\Phi^{4}}{4!}+\frac{1}{2}\xi R\Phi^{2}\right]\qquad\qquad(\lambda>0)$ (1)
The corresponding equation of motion for $\Phi$ reads
$\left(\square+m^{2}+\frac{\eta\Phi}{2}+\frac{\lambda\Phi^{2}}{3!}+\xi R\right)\Phi=0$ (2)
We are interested to do quantum field theory in a neighborhood of a given spacetime point. Accordingly, we expand the metric in the Riemann normal coordinate ($y^{\mu}$) around it [79]
$g_{\mu\nu}=\eta_{\mu\nu}-\frac{1}{3}R_{\mu\alpha\nu\beta}y^{\alpha}y^{\beta}-\frac{1}{6}R_{\mu\alpha\nu\beta;\gamma}y^{\alpha}y^{\beta}y^{\gamma}+\left(-\frac{1}{20}R_{\mu\alpha\nu\beta;\gamma\delta}+\frac{2}{45}R_{\alpha\mu\beta\lambda}R^{\lambda}{}_{\gamma\nu\delta}\right)y^{\alpha}y^{\beta}y^{\gamma}y^{\delta}\cdots$
$|g|=1-\frac{1}{3}R_{\alpha\beta}y^{\alpha}y^{\beta}-\frac{1}{6}R_{\alpha\beta;\gamma}y^{\alpha}y^{\beta}y^{\gamma}+\left(\frac{1}{18}R_{\alpha\beta}R_{\gamma\delta}-\frac{1}{90}R^{\lambda}{}_{\alpha\beta}\,^{\kappa}R_{\lambda\gamma\delta\kappa}-\frac{1}{20}R_{\alpha\beta;\gamma\delta}\right)y^{\alpha}y^{\beta}y^{\gamma}y^{\delta}\cdots$ (3)
The curvature terms appearing above serve as expansion coefficients, evaluated at the origin of the normal coordinate, where $y=0$. The above expansion only holds good if we are interested in a length scale small compared to the characteristic scale associated with the spacetime $\sim R^{-1/2}$, over which the variation of the curvature is effective.
Let $G(x,x^{\prime})$ be the Feynman propagator in this spacetime. For computational convenience, one defines $\bar{G}(x,x^{\prime})$ as
$G(x,x^{\prime})=g^{-{\frac{1}{4}}}(x)\bar{G}(x,x^{\prime})g^{-{\frac{1}{4}}}(x^{\prime})$ (4)
One can next use the local flatness of the spacetime, 3, in order to define a local momentum space, leading to a Fourier decomposition [4, 34]
$\bar{G}(x,x^{\prime})=\int\frac{d^{d}x}{(2\pi)^{d}}e^{ik\cdot y}\bar{G}(k)$ (5)
We now expand $\bar{G}(k)$ as
$\bar{G}(k)=\bar{G_{0}}(k)+\bar{G_{1}}(k)+\bar{G_{2}}(k)+\bar{G_{3}}(k)+\cdots$ (6)
where $\bar{G_{i}}(k)$ contains the $i^{\rm th}$ derivative of the metric at the centre of the Riemann normal coordinate, $y=0$.
We have at the leading order
$\bar{G_{0}}(k)=\frac{1}{k^{2}+m^{2}}$ (7)
Since there is no first order derivative of the metric surviving at $y=0$, we have $\bar{G_{1}}(k)=0$. One also computes
$\bar{G_{2}}(k)=\frac{\left(\frac{1}{6}-\xi\right)R}{(k^{2}+m^{2})^{2}},\qquad\qquad\bar{G_{3}}(k)=\frac{i\left(\frac{1}{6}-\xi\right)(\nabla_{\alpha}R)}{k^{2}+m^{2}}\partial^{\alpha}_{k}\frac{1}{k^{2}+m^{2}}$
$\bar{G_{4}}(k)=\frac{\left(\frac{1}{6}-\xi\right)^{2}{R^{2}}}{(k^{2}+m^{2})^{3}}+\frac{a_{\alpha\beta}}{k^{2}+m^{2}}\partial^{\alpha}_{k}\partial^{\beta}_{k}\frac{1}{k^{2}+m^{2}}$ (8)
where we have abbreviated
$a_{\alpha\beta}=\frac{1}{2}\left(\xi-\frac{1}{6}\right)\nabla_{\alpha}\nabla_{\beta}R+\frac{1}{120}\nabla_{\alpha}\nabla_{\beta}R-\frac{1}{40}\square R_{\alpha\beta}+\frac{1}{30}R_{\alpha}\,^{\lambda}R_{\lambda\beta}-\frac{1}{60}{R^{\rho}\,_{\alpha}}^{\lambda}\,_{\beta}R_{\rho\lambda}-\frac{1}{60}R_{\lambda\mu\rho\alpha}R^{\lambda\mu\rho}{}_{\beta}$ (9)
We shall restrict our calculation to the quadratic order of the curvature only. Up to this order we collect all the terms to write
$\bar{G}(k)=\frac{1}{k^{2}+m^{2}}+\frac{\left(\frac{1}{6}-\xi\right)R}{(k^{2}+m^{2})^{2}}+\frac{i}{2}{\left(\frac{1}{6}-\xi\right)(\nabla_{\alpha}R}){\partial^{\alpha}_{k}}(k^{2}+m^{2})^{-2}+\frac{1}{3}a_{\alpha\beta}\partial^{\alpha}_{k}\partial^{\beta}_{k}(k^{2}+m^{2})^{-2}$
$+\left[\left(\frac{1}{6}-\xi\right)^{2}{R^{2}}-\frac{2}{3}a^{\lambda}\,_{\lambda}\right]\frac{1}{(k^{2}+m^{2})^{3}}+{\cal O}(R^{3})$ (10)
The above equation will serve as the key ingredient for our computations appearing below.
3 One loop effective potential at zero temperature
The computation of the divergent part and subsequent renormalisation of the one loop effective action in a curved spacetime can be seen in, e.g. [4] and references therein. In this section we wish to compute below the finite part of the effective potential at zero temperature after renormalisation, up to quadratic order in the curvature. Although the result appearing in this section can easily be found by directly using the standard ‘$\ln{\rm det}$’ formula for the one loop effective action, we wish to briefly sketch below the use of the local momentum space, for the sake of completeness.
We make the usual decomposition for the field in 1 in terms of the classical background ($\phi$) and fluctuations ($\delta\phi$), $\Phi=\phi+\delta\phi$. Due to the background field the rest mass term appearing in 10 gets modified as, $m^{2}\to m^{2}_{1}$ (say) where
$m^{2}_{1}=m^{2}+\frac{1}{2}\lambda\phi^{2}+\eta\phi$
Figure 1: The one loop diagrams for quartic plus cubic self interactions. The above three rows respectively contain quartic, cubic and mixed interaction vertices. Each row contains infinite number of one loop diagrams, found by inserting vertices on the vacuum loop. The vertices correspond to the background field, and no momenta flow across them.
1 represents the diagrammatic expansion of the one loop effective action. The same is basically the sum of all one loop diagrams found via vertex insertions shown in this figure. In the Minkowski spacetime, the one loop effective potential reads
$V_{\rm{eff}}^{1-{\rm loop}}=\frac{\hbar}{2}\int\frac{d^{d}k}{(2\pi)^{d}}\ln(k^{2}+m^{2}_{1})=-\frac{\hbar}{2}\int\frac{d^{d}k}{(2\pi)^{d}}\ln G(k,m_{1})$ (11)
where $G(k,m_{1})$ is the propagator of the scalar in the flat spacetime with effective mass $m_{1}$. We now replace the flat space propagator by that of the curved spacetime, 10, to find
$V_{\rm{eff}}^{1-{\rm loop}}=-\frac{\hbar}{2}\int\frac{d^{d}k}{(2\pi)^{d}}\ln\left[\frac{1}{(k^{2}+m^{2}_{1})}+\frac{(\frac{1}{6}-\xi)R}{(k^{2}+m_{1}^{2})^{2}}+\frac{2f_{2}}{(k^{2}+m_{1}^{2})^{3}}\right]$
$=\frac{\hbar}{2}\int\frac{d^{d}k}{(2\pi)^{d}}\left[\ln(k^{2}+m^{2}_{1})-\frac{(\frac{1}{6}-\xi)R}{(k^{2}+m_{1}^{2})}-\frac{2f_{2}}{(k^{2}+m_{1}^{2})^{2}}+\frac{1}{2}\frac{(\frac{1}{6}-\xi)^{2}R^{2}}{(k^{2}+m_{1}^{2})^{2}}\right]+{\cal O}(R^{3})$ (12)
where the coefficient $f_{2}$ is quadratic in curvature,
$f_{2}=\frac{1}{2}\left(\frac{1}{6}-\xi\right)^{2}{R^{2}}-\frac{1}{3}a^{\lambda}\,_{\lambda}$
and $a_{\mu\nu}$ is given by 9. The first term on the right hand side of 12 is the flat space term, whereas the rest are the leading curvature corrections. For $d=4-\epsilon$, the standard momentum integrals appearing above are given by
$\int\frac{d^{d}k}{(2\pi)^{d}}\ln(k^{2}+m_{1}^{2})=\frac{m^{4}_{1}}{(4\pi)^{2}}\left[-\frac{\mu^{-\epsilon}}{\epsilon}+\frac{1}{2}\ln\frac{m_{1}^{2}}{4\pi\mu^{2}}-\frac{1}{2}-\frac{1}{2}\psi(2)\right]+{\cal O}(\epsilon)$
$-\left(\frac{1}{6}-\xi\right)R\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{(k^{2}+m_{1}^{2})}=-\left(\frac{1}{6}-\xi\right)\frac{Rm^{2}_{1}}{(4\pi)^{2}}\left[-\frac{2\mu^{-\epsilon}}{\epsilon}+\ln\frac{m_{1}^{2}}{4\pi\mu^{2}}-\psi(2)\right]+{\cal O}(\epsilon)$
$\left(\frac{1}{2}{\left(\frac{1}{6}-\xi\right)^{2}R^{2}}-2f_{2}\right)\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{(k^{2}+m_{1}^{2})^{2}}=\frac{{\left(\frac{1}{6}-\xi\right)^{2}R^{2}-4f_{2}}}{2(4\pi)^{2}}\left[\frac{2\mu^{-\epsilon}}{\epsilon}+\psi(1)-\ln\frac{m^{2}_{1}}{4\pi\mu^{2}}\right]+{\cal O}(\epsilon)$ (13)
where $\psi(z)=\Gamma^{\prime}(z)/\Gamma(z)$ is the digamma function and $\mu$ is a renormalisation scale. Substituting the three above integrals into 12 yields the (unrenormalised) one loop effective action. The renormalisation of the effective potential or the action can be seen in [4]. The divergent part of the effective potential reads (with $\hbar=1$)
$V_{\rm{eff,\ div.}}^{1-{\rm loop}}=-\frac{\mu^{-\epsilon}}{16\pi^{2}\epsilon}\left[\left(\frac{1}{72}+\frac{\xi^{2}}{2}-\frac{\xi}{6}\right)R^{2}+\frac{1}{90}\left(R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}-R_{\mu\nu}R^{\mu\nu}\right)+\frac{m^{4}}{2}+m^{2}\eta\phi+\left(\xi-\frac{1}{6}\right)m^{2}R\right.$
$\left.+\left(\xi-\frac{1}{6}\right)\eta R\phi+\frac{1}{2}\left(m^{2}\lambda+\eta^{2}\right)\phi^{2}+\frac{\lambda^{2}\phi^{4}}{8}+\frac{\eta\lambda\phi^{3}}{2}+\frac{\lambda}{2}\left(\xi-\frac{1}{6}\right)R\phi^{2}-\frac{1}{3}\left(\xi-\frac{1}{5}\right)\square R\right]$ (14)
We may ignore the last term as it is a total divergence. In order to renormalise the divergence associated with the pure curvature terms, we take the gravitational counterterms in the Lagrangian density
$\delta{\cal L}_{\rm grav}=\left[\delta\Lambda+\delta\kappa R+\delta\alpha_{1}R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}+\delta\alpha_{2}R_{\mu\nu}R^{\mu\nu}+\delta\alpha_{3}R^{2}\right],$ (15)
with the choice
$\delta\Lambda=\frac{\mu^{-\epsilon}m^{4}}{32\pi^{2}\epsilon},\qquad\delta\kappa=\frac{\mu^{-\epsilon}m^{2}\left(\xi-\frac{1}{6}\right)}{16\pi^{2}\epsilon},\qquad\delta\alpha_{1}=\frac{\mu^{-\epsilon}}{16\pi^{2}\epsilon}\frac{1}{90}$
$\delta\alpha_{2}=-\frac{\mu^{-\epsilon}}{16\pi^{2}\epsilon}\frac{1}{90},\qquad\delta\alpha_{3}=\frac{\mu^{-\epsilon}}{32\pi^{2}\epsilon}\left(\frac{1}{36}+\xi^{2}-\frac{\xi}{3}\right)\qquad$ (16)
The remaining divergence of 14 is absorbed by a matter counterterm Lagrangian
$\delta{\cal L}_{\rm matt}=\left[\frac{1}{2}\delta m^{2}\phi^{2}+\frac{1}{2}\delta\xi R\phi^{2}+\frac{\delta\eta\phi^{3}}{3!}+\frac{\delta\lambda\phi^{4}}{4!}+\delta\tau\phi+\delta\gamma R\phi\right],$ (17)
with the choice
$\delta m^{2}=-\frac{\mu^{-\epsilon}\left(m^{2}\lambda+\eta^{2}\right)}{16\pi^{2}\epsilon},\qquad\delta\xi=-\frac{\mu^{-\epsilon}\lambda\left(\xi-\frac{1}{6}\right)}{16\pi^{2}\epsilon},\qquad\delta\tau=-\frac{\mu^{-\epsilon}m^{2}\eta}{16\pi^{2}\epsilon}$
$\delta\gamma=-\frac{\mu^{-\epsilon}\eta\left(\xi-\frac{1}{6}\right)}{16\pi^{2}\epsilon},\qquad\delta\eta=-\frac{3\mu^{-\epsilon}\eta\lambda}{16\pi^{2}\epsilon},\qquad\delta\lambda=-\frac{3\mu^{-\epsilon}\lambda^{2}}{16\pi^{2}\epsilon}$ (18)
We further define some finite counterterms to absorb terms which are independent of $\phi$ but otherwise finite,
$\delta\Lambda_{\rm fin.}=\frac{m^{4}}{64\pi^{2}}\left(\frac{1}{2}+\psi(2)\right),\qquad\delta\kappa_{\rm fin.}=\frac{m^{2}\psi(2)\left(\xi-\frac{1}{6}\right)}{32\pi^{2}}$
$\delta\alpha_{1,{\rm fin.}}=\frac{1}{5760\pi^{2}},\qquad\delta\alpha_{2,{\rm fin.}}=\frac{1}{5760\pi^{2}},\qquad\delta\alpha_{3,{\rm fin.}}=\frac{\left(\xi-\frac{1}{6}\right)^{2}}{4(4\pi)^{2}},$ (19)
to finally have the renormalised result
$V_{\rm{eff,Ren}}^{1-{\rm loop}}=\frac{1}{2}(m^{2}+\xi R)\phi^{2}+\frac{\lambda\phi^{4}}{4!}+\frac{\eta\phi^{3}}{3!}+\frac{1}{2(4\pi)^{2}}\left[\frac{m^{4}_{1}}{2}\left(\ln\frac{m_{1}^{2}}{4\pi\mu^{2}}-\frac{1}{2}-\psi(2)\right)+\frac{m^{4}}{2}\left(\frac{1}{2}+\psi(2)\right)\right.$
$\left.+\left(\frac{1}{6}-\xi\right)R\left(\psi(2)(m^{2}_{1}-m^{2})-m_{1}^{2}\ln\frac{m_{1}^{2}}{4\pi\mu^{2}}\right)-\left(\frac{1}{2}\left(\frac{1}{6}-\xi\right)^{2}R^{2}-2f_{1}\right)\ln\frac{m^{2}_{1}}{4\pi\mu^{2}}\right]$ (20)
where $m_{1}^{2}=m^{2}+\frac{1}{2}\lambda\phi^{2}+\eta\phi$, and we have abbreviated
$f_{1}=\frac{1}{2}\left(\frac{1}{6}-\xi\right)^{2}R^{2}-\frac{1}{3}a^{\lambda}\,_{\lambda}=\frac{1}{2}\left(\frac{1}{6}-\xi\right)^{2}{R^{2}}+\frac{1}{180}R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}-\frac{1}{180}R_{\mu\nu}R^{\mu\nu}$
Let us now take the flat spacetime limit of 20, for a massless and quartic self interction. Taking $\psi(2)=0.422785$, we obtain the Coleman-Weinberg effective potential [5]
$V_{\rm{eff}}^{\rm flat}$ $=\frac{\lambda\phi^{4}}{4!}+\frac{\lambda^{2}\phi^{4}}{256\pi^{2}}\left[\ln\frac{\lambda\phi^{2}}{\mu^{2}}-\ln(8\pi)-\frac{1}{2}-0.422785\right]$ (21)
$=\frac{\lambda\phi^{4}}{4!}+\frac{\lambda^{2}\phi^{4}}{256\pi^{2}}\left[\ln\frac{\lambda\phi^{2}}{\mu^{2}}-4.146\right]$
We note a slight mismatch between the numerical factor appearing above ($4.146$), with that of [5], which is $4.166$.
4 The case at finite temparature
We next wish to generalise 12 for finite temperature, $T=\beta^{-1}$. In order to obtain the pure thermal contribution, we recall that one needs replacing the Euclidean momentum $k_{0}$ by discrete $\omega_{n}$ where $\omega_{n}=2\pi n/\beta$ for bosons, i.e. [20, 67]
$k^{2}+m_{1}^{2}\to\vec{k}^{2}+\left(\frac{2\pi n}{\beta}\right)^{2}+m_{1}^{2},\qquad\qquad n=0,\pm 1,\pm 2,\pm 3,\cdots$ (22)
This leads to the replacements in 12
$\int\frac{d^{d}k}{(2\pi)^{d}}\ln(k^{2}+m_{1}^{2})\longrightarrow\sum_{n=-\infty}^{\infty}\frac{1}{\beta}\int\frac{d^{d-1}\vec{k}}{(2\pi)^{d-1}}\ln\left[\vec{k}^{2}+\left(\frac{2\pi n}{\beta}\right)^{2}+m_{1}^{2}\right]$
$\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{(k^{2}+m_{1}^{2})^{s}}\longrightarrow\sum_{n=-\infty}^{\infty}\frac{1}{\beta}\int\frac{d^{d-1}\vec{k}}{(2\pi)^{d-1}}\frac{1}{\left[\vec{k}^{2}+\left(\frac{2\pi n}{\beta}\right)^{2}+m_{1}^{2}\right]^{s}}$ (23)
Before we proceed for the effective action, let us see the effect of the curvature on the self energy in a thermal field theory, in order to further motivate the problem we are interested in. For a scalar field with mass $m$ and quartic self interaction, the renormalised one loop self energy reads [25]
$\Sigma=12\lambda\int\frac{d^{3}\vec{k}}{(2\pi)^{3}}\frac{1}{\left(\vec{k}^{2}+m^{2}\right)^{\frac{1}{2}}}\frac{1}{e^{\beta\sqrt{\vec{k}^{2}+m^{2}}}-1}$ (24)
Assuming the scalar to be light, the first two leading order terms read
$\Sigma=\lambda T^{2}\left(1-\frac{3m}{\pi T}\right)$ (25)
This shows generation of a scalar mass due to the finite temperature. Let us now ‘turn on’ the gravity and for simplicity assume the spacetime to be weakly curved. Keeping only up to the ${\cal O}(R)$ term, we have from 10
${\bar{G}}(k)=\left[1-\left(\frac{1}{6}-\xi\right)R\frac{\partial}{\partial m^{2}}\right]\frac{1}{k^{2}+m^{2}}$ (26)
Promoting now the propagator to thermal case as stated above, we have the renormalised one loop self energy
$\Sigma=\lambda T^{2}-\frac{3\lambda mT}{\pi}-\left(\frac{1}{6}-\xi\right)R\frac{\partial}{\partial m^{2}}\left(\lambda T^{2}-\frac{3\lambda mT}{\pi}\right)=\lambda T^{2}-\frac{3\lambda mT}{\pi}+\left(\frac{1}{6}-\xi\right)\frac{3\lambda RT}{2\pi m}$ (27)
The third term of the final result exhibits mass generation in a thermal field theory in a weakly curved spacetime. We note that if the scalar is extremely light, the curvature correction term becomes very large, making the result untrustworthy. It seems that one needs to develop some non-perturbative resummation technique in order to tackle such issue. Nevertheless, the above result indeed indicates that the spacetime curvature can have interesting effects on thermal fields in a curved spacetime.
Let us now come to the computation of the effective potential. In 23, we may break the summation by separating the zero mode as
$\sum_{n=-\infty}^{\infty}=(n=0)+2\sum_{n=1}^{\infty}$ (28)
We shall essentially work in the high temperature approximation $(m\beta\ll 1)$ and hence will make a power series expansion in $\beta$ for relevant quantities, and will retain terms up to ${\cal O}(\beta^{6})$ in the following.
Now, the pure thermal part of the first integral of 13 becomes at finite temperature
$\sum_{n=1}^{\infty}\frac{1}{\beta}\int\frac{d^{d-1}\vec{k}}{(2\pi)^{d-1}}\ln\left[\vec{k}^{2}+\left(\frac{2\pi n}{\beta}\right)^{2}+m_{1}^{2}\right]=\sum_{n=1}^{\infty}\frac{1}{\beta}\frac{2}{(d-1)(4\pi)^{\frac{d-1}{2}}}\left[\left(\frac{2\pi n}{\beta}\right)^{2}+m_{1}^{2}\right]^{\frac{d-1}{2}}\Gamma\left(\frac{3-d}{2}\right)$ (29)
We have,
$\sum_{n=1}^{\infty}\left[\left(\frac{2\pi n}{\beta}\right)^{2}+m_{1}^{2}\right]^{\frac{d-1}{2}}=\sum_{n=1}^{\infty}m_{1}^{d-1}\left(n\alpha\right)^{d-1}\left[1+\frac{1}{(n\alpha)^{2}}\right]^{\frac{d-1}{2}}$ (30)
where we have defined $\alpha=2\pi/(m_{1}\beta)$. Expanding for a small $\beta$, the above expression equals
$m_{1}^{d-1}\alpha^{-\epsilon}\left[\alpha^{3}\zeta(-3)+\frac{3}{2}\alpha\zeta(-1)+\frac{3}{8}\alpha^{-1}\zeta(1+\epsilon)-\frac{1}{16}\alpha^{-3}\zeta(3)+\frac{1}{128}\alpha^{-5}\zeta(5)-\frac{5}{3072}\alpha^{-7}\zeta(3)+{\cal O}(\beta^{8})\right]$ (31)
where $\zeta(n)=\sum_{j=1}^{\infty}j^{-n}$ is the Riemann zeta function. Using the properties of this function pertaining to its regularisation, 29 becomes
$\sum_{n=1}^{\infty}\frac{1}{\beta}\int\frac{d^{d-1}\vec{k}}{(2\pi)^{d-1}}\ln\left[\vec{k}^{2}+\left(\frac{2\pi n}{\beta}\right)^{2}+m_{1}^{2}\right]=-\frac{1}{90}\frac{\pi^{2}}{\beta^{4}}+\frac{1}{24}\frac{m_{1}^{2}}{\beta^{2}}-\frac{m_{1}^{3}}{12\pi\beta}-\frac{m^{4}_{1}\ln{m^{2}_{1}\beta^{2}}}{64\pi^{2}}+\frac{m^{4}_{1}\left(\psi(1)+\ln{4\pi^{2}}\right)}{64\pi^{2}}$
$+\frac{m^{6}_{1}\beta^{2}\zeta(3)}{768\pi^{4}}-\frac{m^{8}_{1}\beta^{4}\zeta(5)}{24576\pi^{6}}+\frac{m^{10}_{1}\beta^{6}\zeta(7)}{196608\pi^{8}}$ (32)
Likewise we have
$-\sum_{n=1}^{\infty}\left(\frac{1}{6}-\xi\right)R\int\frac{d^{d-1}\vec{k}}{(2\pi)^{d-1}}\frac{1}{\left[\vec{k}^{2}+\left(\frac{2\pi n}{\beta}\right)^{2}+m_{1}^{2}\right]}=-\left(\frac{1}{6}-\xi\right)R\left[\frac{1}{24\beta^{2}}-\frac{m_{1}}{8\pi\beta}+\frac{m_{1}^{2}}{16\pi^{2}}\left(\psi(1)+\ln{\frac{2\pi}{m_{1}\beta}}\right)\right.$
$\left.+\frac{m_{1}^{4}\beta^{2}}{256\pi^{4}}\zeta(3)-\frac{m_{1}^{6}\beta^{4}}{2048\pi^{6}}\zeta(5)+\frac{15m_{1}^{8}\beta^{6}}{24576\pi^{8}}\zeta(7)\right]$ (33)
and
$\sum_{n=1}^{\infty}\left(\frac{1}{2}\left(\frac{1}{6}-\xi\right)^{2}R^{2}-2f_{1}\right)\int\frac{d^{d-1}\vec{k}}{(2\pi)^{d-1}}\frac{1}{\left[\vec{k}^{2}+\left(\frac{2\pi n}{\beta}\right)^{2}+m_{1}^{2}\right]^{2}}=\left(\frac{1}{2}\left(\frac{1}{6}-\xi\right)^{2}R^{2}-2f_{1}\right)\left[\frac{1}{16\pi\beta m_{1}}\right.$
$\left.-\frac{1}{16\pi^{2}}\left(\psi(1)+\ln{\frac{2\pi}{m_{1}\beta}}\right)-\frac{m_{1}^{2}\beta^{2}}{128\pi^{4}}\zeta(3)+\frac{3m_{1}^{4}\beta^{4}}{2048\pi^{6}}\zeta(5)-\frac{15m_{1}^{6}\beta^{6}}{3072\pi^{8}}\zeta(7)\right]$ (34)
An alternative and more rigorous derivation of the above integrals in the flat spacetime by breaking the Feynman propagator into temperature independent and dependent parts can be seen in [7]. We also note that in the three integrals appearing above there are a couple of temperature independent parts, such as the fifth term appearing on the right hand side of 32, even though 32, 33, 34 are pure thermal contributions. Such $\beta$ independent parts originate from the regularisation of the $\zeta$-function.
Putting things together now, we obtain the total renormalised one loop effective potential at the high temperature limit
$V_{\rm{eff},\beta}^{1-{\rm loop}}(\beta^{-1}\gg 1)=\frac{1}{2}(m^{2}+\xi R)\phi^{2}+\frac{\lambda\phi^{4}}{4!}+\frac{\beta\phi^{3}}{3!}+V_{\rm{eff}}^{1-{\rm loop}}(\beta^{-1}=0)+\left[-\frac{\pi^{2}}{90\beta^{4}}+\frac{m_{1}^{2}}{24\beta^{2}}-\frac{m_{1}^{3}}{12\pi\beta}\right.$
$\left.-\frac{m^{4}_{1}\ln{m^{2}_{1}\beta^{2}}}{64\pi^{2}}+\frac{m^{4}_{1}\left(\psi(1)+\ln{4\pi}\right)}{64\pi^{2}}+\frac{m^{6}_{1}\beta^{2}\zeta(3)}{768\pi^{4}}-\frac{m^{8}_{1}\beta^{4}\zeta(5)}{24576\pi^{6}}+\frac{m^{10}_{1}\beta^{6}\zeta(7)}{196608\pi^{8}}\right.$
$\left.-\left(\frac{1}{6}-\xi\right)R\left(\frac{1}{24\beta^{2}}-\frac{m_{1}}{8\pi\beta}+\frac{m_{1}^{2}}{16\pi^{2}}\left(\psi(1)+\ln\frac{2\pi}{m_{1}\beta}\right)+\frac{m_{1}^{4}\beta^{2}}{256\pi^{4}}\zeta(3)-\frac{m_{1}^{6}\beta^{4}}{6144\pi^{6}}\zeta(5)+\frac{5m_{1}^{8}\beta^{6}}{196608\pi^{8}}\zeta(7)\right)\right.$
$\left.+\left(\frac{1}{2}\left(\frac{1}{6}-\xi\right)^{2}R^{2}-2f_{1}\right)\left(\frac{1}{16\pi\beta m_{1}}-\frac{1}{16\pi^{2}}\left(\psi(1)+\ln\frac{2\pi}{m_{1}\beta}\right)-\frac{m_{1}^{2}\beta^{2}}{128\pi^{4}}\zeta(3)+\frac{m_{1}^{4}\beta^{4}}{2048\pi^{6}}\zeta(5)-\frac{5m_{1}^{6}\beta^{6}}{49152\pi^{8}}\zeta(7)\right)+{\cal O}(\beta^{7})\right]$
where $m_{1}^{2}=m^{2}+\lambda\phi^{2}/2+\eta\phi$, as earlier and $V_{\rm{eff}}^{1-{\rm loop}}(\beta^{-1}=0)$ is the zero temperature effective potential given by 20. Note that there is no temperature dependent divergence present in the above expression. This shows that renormalisation counterterms are independent of $\beta$, as is expected [25].
As we have mentioned earlier, the above one loop effective potential was first derived in [67] for $\eta=0$. However, our notation is different compared to this work. We also note from [67] that if we set the arbitrary mass scale $\mu=\beta^{-1}$ in 20, all the logarithmic terms appearing in the finite temperature part of the effective potential precisely cancel with that of the zero temperature. However, we wish to keep $\mu$ to be a free parameter in this work. Such terms will contribute non-trivially to the feature of the effective potential and hence to the spontaneous symmetry breaking phenomenon.
Although we are interested in the high temperature limit of the effective action, we wish to briefly discuss the low temperature limit of the same below. For $\beta^{-1}\to 0$, we make the expansion
$\left[\left(\frac{2\pi n}{\beta}\right)^{2}+m_{1}^{2}\right]^{\frac{d-1}{2}}=m_{1}^{3}\left[1+\frac{3}{2}\left(\frac{2\pi n}{m_{1}\beta}\right)^{2}+\frac{3}{8}\left(\frac{2\pi n}{m_{1}^{2}\beta}\right)^{4}+\cdots\right]$ (36)
Putting this expression back into 29, we obtain
$\sum_{n=1}^{\infty}\frac{1}{\beta}\int\frac{d^{d-1}\vec{k}}{(2\pi)^{d-1}}\ln\left[\vec{k}^{2}+\left(\frac{2\pi n}{\beta}\right)^{2}+m_{1}^{2}\right]_{\beta^{-1}\to 0}\approx\sum_{n=1}^{\infty}\frac{1}{\beta}\frac{2m_{1}^{3}}{(d-1)(4\pi)^{\frac{d-1}{2}}}\left(1+\frac{3}{2}\left(\frac{2\pi n}{m_{1}\beta}\right)^{2}+\frac{3}{8}\left(\frac{2\pi n}{m_{1}\beta}\right)^{4}\right)\Gamma\left(\frac{3-d}{2}\right)$
The other relevant integrals can be approximated in a likewise manner. Using now Ramanujan’s result, $\zeta{(-2k)}=0$ ($k=\ {\rm integer}$), we obtain after a little bit of algebra
$V_{\rm{eff},\beta^{-1}\ll 1}=\frac{1}{2}m^{2}\phi^{2}+\frac{\eta\phi^{3}}{3!}+\frac{\lambda\phi^{4}}{4!}+\frac{\xi R\phi^{2}}{2}+V_{\rm{eff}}(\beta^{-1}=0)$
$+\frac{3}{2}\left[-\frac{m_{1}^{3}}{3\pi\beta}+\left(\frac{1}{6}-\xi\right)\frac{m_{1}R}{2\pi\beta}+\left(-f_{1}+\frac{1}{4}\left(\frac{1}{6}-\xi\right)^{2}R^{2}\right)\frac{1}{\pi m_{1}\beta}\right]$ (38)
Setting $\beta^{-1}\to 0$ is possible in 38 whereas it is not possible in 35.
We further wish to consider two loop computations for the effective potential. However, before we do that, let us discuss the issue of spontaneous symmetry breaking and phase transition with the one loop high temperature effective potential, 35, we have found. We also note in 35 that for $\eta\neq 0$, and since both $\eta$ and $\phi$ can be positive or negative, the expression becomes invalid for $m_{1}^{2}\to 0$, owing to the $\ln m_{1}^{2}$ term. In other words, the effective potential will be valid in the parameter or the field space region where $m_{1}^{2}>0$. Owing to the fact that the background field $\phi$ is real, it is easy to see that this condition implies,
$\left(\frac{\eta}{m}\right)^{2}<2\lambda$ (39)
We shall keep this in mind in the following computations whenever relevant.
5 Spontaneous symmetry breaking and phase transition at zero temperature
The chief motivation of this section is to see the effect of the spacetime curvature in the phase transition or spontaneous symmetry breaking phenomenon, at zero temperature. As we have already mentioned, this topic is old. In flat spacetime, one of the earliest thermal field theory calculations in this context can be seen in the seminal work [7]. At zero temperature and with spacetime curvature some such computations can be seen in e.g. [4, 33, 38] and references therein. In particular, [33, 38] use weak or linear in curvature approximation and renormalisation group techniques. However as we have also mentioned earlier, such results cannot be applicable to Ricci flat spacetimes.
Let us now make the field and the cubic coupling constant dimensionless by scaling them with respect to its rest mass, which we assume to be non-vanishing, i.e., $\bar{\phi}=\phi/m$ and $\bar{\eta}=\eta/m$. The tree level asymmetric self interaction potential has been plotted in 2.
Figure 2: Plot of the tree level dimensionless self interaction potential, $V(\bar{\phi})=\bar{\eta}\bar{\phi}^{3}/3!+\lambda\bar{\phi}^{4}/4!$. We have taken $\lambda=0.1$ and $\bar{\eta}=0.447$.
Let us now consider the zero temperature effective potential in the absence of curvature, 20 ($R=0=f_{1}$), in the dimensionless form
$\frac{V_{\rm{eff}}}{m^{4}}=\frac{\bar{\phi}^{2}}{2}+\frac{\bar{\eta}\bar{\phi}^{3}}{3!}+\frac{\lambda\bar{\phi}^{4}}{4!}+\frac{1}{4(4\pi)^{2}}\left[\left(2\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)+\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{2}\right)\left(\ln\frac{1+\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}}{\frac{4\pi\mu^{2}}{m^{2}}}-\frac{1}{2}-\psi(2)\right)\right.$
$\left.+\ln\left(1+\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)\right]$ (40)
Putting in now some customary numerical values, $m\sim 0.00051{\rm GeV}$ and $\mu\sim 3\times 10^{-25}{\rm GeV}$, which is appropriate for the large scale physics we are interested in, e.g. [80] and references therein, we have plotted the above zero temperature flat space effective potential in 3. We note that since the scalar we have taken has non-vanishing rest mass, there can be no spontaneous symmetry breaking for the quartic self interaction. However, in the presence of spacetime curvature, the scenario can change, as we wish to show below.
(a) $\eta\phi^{3}+\lambda\phi^{4}$ interaction, $\lambda$=0.1
(b) $\lambda\phi^{4}$ interaction
Figure 3: The one loop effective potential, 40, with vanishing temperature in flat spacetime. Note that for $\bar{\eta}=0$, no spontaneous symmetry breaking is indicated, as we have taken the rest mass of the scalar to be non-vanishing.
We wish to consider two physically well motivated spacetimes below, namely, the Schwarzschild and the de Sitter spacetime.
The Schwarzschild spacetime:
For the Schwarzschild spacetime,
$ds^{2}=\left(1-\frac{2MG}{r}\right)dt^{2}+\left(1-\frac{2MG}{r}\right)^{-1}dr^{2}+r^{2}d\Omega^{2},$
we must set $R=0=R_{\mu\nu}$, and
$R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}=\frac{48G^{2}M^{2}}{r^{6}},$
in 20, so that we have around $r\sim 2GM$
$\frac{V_{\rm{eff}}}{m^{4}}=\frac{\bar{\phi}^{2}}{2}+\frac{\bar{\eta}\bar{\phi}^{3}}{3!}+\frac{\lambda\bar{\phi}^{4}}{4!}+\frac{1}{4(4\pi)^{2}}\left[\left(2\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)+\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{2}\right)\left(\ln\frac{1+\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}}{\frac{4\pi\mu^{2}}{m^{2}}}-\frac{1}{2}-\psi(2)\right)\right.$
$\left.+\ln\left(1+\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)\right]+\frac{0.0016}{m^{4}G^{4}M^{4}}\ln\left(1+\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)$ (41)
Note that the effective potential diverges as $M\to 0$. This corresponds to the fact that the value of $R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}$ increases on the horizon with decreasing mass of the black hole. On the other hand, the result approaches the flat spacetime limit, 40, for large value of the black hole mass. Taking now for example, $GM\sim 10^{2}{\rm m}$, we have plotted 41 in 4. We note in particular for the first of 4, that for increasing cubic coupling, the effective potential develops a deep. This originates from the logarithm contained in the last term on the right hand side of 41, which may contribute a large negative value. Note also that for a given $\lambda$ value, we cannot arbitrarily go on increasing the cubic coupling, owing to the discussion appearing at the end of the preceding Section, 39.
(a) $\eta\phi^{3}+\lambda\phi^{4}$ interaction, $\lambda=0.1$
(b) $\lambda\phi^{4}$ interaction
Figure 4: The one loop effective potential, 41, at vanishing temperature in the Schwarzschild spacetime. See main text for discussion.
Let us also consider in this context the Kerr spacetime,
$ds^{2}=-\frac{r^{2}-2GMr+a^{2}}{r^{2}+a^{2}\cos^{2}\theta}dt^{2}-\frac{4aMGr\sin^{2}\theta}{r^{2}+a^{2}\cos^{2}\theta}dtd\phi+\frac{r^{2}+a^{2}\cos^{2}\theta}{r^{2}-2MGr+a^{2}}dr^{2}+\left(r^{2}+a^{2}\cos^{2}\theta\right)d\theta^{2}$
$+\left(r^{2}+a^{2}+\frac{2a^{2}GMr\sin^{2}\theta}{r^{2}+a^{2}\cos^{2}\theta}\right)\sin^{2}\theta d\phi^{2}$ (42)
which is also Ricci flat, and the Kretschmann scalar reads
$R_{\mu\nu\lambda\rho}R^{\mu\nu\lambda\rho}=\frac{48G^{2}M^{2}\left(r^{6}-15a^{2}r^{4}\cos^{2}\theta+15a^{4}r^{2}\cos^{2}\theta-a^{6}\cos^{6}\theta\right)}{\left(r^{2}+a^{2}\cos^{2}\theta\right)^{6}}$
Comparing this with the Schwarzschild spacetime, we see that there is no qualitative change at all due to spacetime rotation on the equatorial plane, $\theta=\pi/2$. The effect of the spacetime rotation is maximum on the axis ($\theta=0,\pi$). We note that if we take the extremal limit ($a\to GM$), the Kretschmann scalar vanishes as we move towards the horizon ($r\to GM$) along the axis. This shows that the one loop effective potential for an extreme Kerr black hole spacetime in the vicinity of the axial points near the horizon coincides with that of the flat spacetime, 40.
The de Sitter spacetime:
Let us now come to the de Sitter spacetime,
$ds^{2}=-dt^{2}+e^{2\sqrt{\Lambda}t/\sqrt{3}}\left(dx^{2}+dy^{2}+dz^{2}\right)$
for which we have $R=4\Lambda$, $R^{\mu\nu}R_{\mu\nu}=4\Lambda^{2}$ and $R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}=8\Lambda^{2}/3$ in 20, $\Lambda$ being the positive cosmological constant. We have the one loop effective potential
$\frac{V_{\rm{eff}}}{m^{4}}=\frac{\bar{\phi}^{2}}{2}+\frac{\bar{\eta}\bar{\phi}^{3}}{3!}+\frac{\lambda\bar{\phi}^{4}}{4!}+\frac{2\xi\Lambda\bar{\phi}^{2}}{m^{2}}$
$+0.0015\times\left[\left(2\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)+\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{2}\right)\left(\ln\frac{1+\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}}{\frac{4\pi\mu^{2}}{m^{2}}}-\frac{1}{2}-\psi(2)\right)+\ln\left(1+\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)\right]$
$-0.012\times\frac{\left(\frac{1}{6}-\xi\right)\Lambda}{m^{2}}\left[\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)\left(\ln\frac{1+\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}}{\frac{4\pi\mu^{2}}{m^{2}}}-\psi(2)\right)+\ln\left(1+\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)\right]$
$+\frac{0.032\Lambda^{2}}{m^{4}}\ln\left(1+\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)\left[\left(\frac{1}{6}-\xi\right)^{2}-0.001\right]$ (43)
We have investigated various features of this potential in 5, 6 and 7.
(a) $\eta\phi^{3}+\lambda\phi^{4}$ interaction, $\lambda=0.1$
(b) $\lambda\phi^{4}$ interaction
Figure 5: The one loop effective potential, 43, at vanishing temperature in the de Sitter spacetime. Comparing the latter with 3(b), 4(b), we see that spontaneous symmetry breaking in de Sitter with a massive scalar with $m^{2}{>}0$ is possible. In both the plots appearing above we have taken $\Lambda\sim 10^{-52}{\rm m}^{-2}$, i.e. its current observed value and $\xi\sim 0.13$. See main text for discussion.
In particular, we note that (with $\bar{\eta}=0$), spontaneous symmetry breaking at one loop is possible in de Sitter even with $m^{2}{>}0$. This has no analogue for flat or the Ricci flat spacetimes we have discussed, 3(b), 4(b). In fact symmetry breaking for a massive scalar in flat spacetime seems only to be possible if the scale $\mu$ is extremely large compared to $m$. In de Sitter, this effect originates from the term appearing in the last line of the right hand side of 43, which, when $\xi$ is $1/6$ or close to it, yields a negative term which can dominate the $\bar{\phi}^{2}/2$ term. Eventually this leads to the symmetry breaking effective potential. We see that this is possible with a $\Lambda$ as tiny as its current observed value. Certainly, such effect may a priori be expected to be much more stronger in the early inflationary universe scenario, where the dark energy density was much higher compared to that of today. Note also from the second of 7 that symmetry breaking remains possible for a very wide range of the renormalisation scale $\mu$. Finally, we note that had we ignored the quadratic curvature terms from the beginning, we would not have seen any such effect. We believe this feature, along with 27, sufficiently motivate us to look into this problem by employing some non-perturbative resummation techniques. Such technique should not only resum the coupling, but must resum some curvature terms as well. We reserve this problem for a future work.
(a) $\lambda\phi^{4}$ interaction
(b) $\lambda\phi^{4}$ interaction
Figure 6: The one loop effective potential, 43, at vanishing temperature in the de Sitter spacetime for different values of the non-minimal coupling $\xi$. Note that for $\xi\sim 0.167$ or $1/6$, the dip or the hill of the effective potential is maximum. This effect originates from the terms appearing on the last line on the right hand side of 43. See main text for discussion.
(a) $\lambda\phi^{4}$ interaction
(b) $\lambda\phi^{4}$ interaction
Figure 7: The one loop effective potential, 43, with vanishing temperature in the de Sitter spacetime for different values of the cosmological constant $\Lambda$, and the mass scale $\mu$. Increasing the value of $\Lambda$ increases the strength of the symmetry breaking effect. Note also that the symmetry breaking feature of the potential remains valid for a very wide range of the renormalisation scale, $\mu$.
We now wish to discuss the above effective potentials in the context of finite temperature field theory.
5.1 Phase transition at finite temperature
$m^{2}{<}0$ case :
To begin with, let us consider the thermal effective potential 35 at zero spacetime curvature with only quartic coupling and at high temperature, e.g. [27],
$V_{\rm{eff},\beta}=\frac{m^{2}\phi^{2}}{2}+\frac{\lambda\phi^{4}}{4!}+\left[-\frac{\pi^{2}}{90\beta^{4}}+\frac{m^{2}+\frac{1}{2}\lambda\phi^{2}}{24\beta^{2}}\right]\qquad(\lambda>0)$ (44)
where we have kept only the first two thermal terms for simplicity. It is well known that for $m^{2}{<}0$ there can be symmetry breaking for the tree level potential, as has been depicted in 8.
Figure 8: Tree level symmetry breaking potential with $m^{2}<0$ and positive quartic self interaction.
Following e.g. [27], we now rewrite 44 as
$V_{\rm{eff},\beta}=\frac{m^{2}\phi^{2}}{2}\left(1-\frac{T^{2}}{T^{2}_{c}}\right)+\frac{\lambda\phi^{4}}{4!}$ (45)
Where, $T_{c}=\sqrt{-\frac{24m^{2}}{\lambda}}$ is called the critical temperature, and we have ignored the two $\phi$-independent terms. Since $m^{2}{<}0$, there is always a symmetry breaking for $T{<}T_{c}$, but for $T\geq T_{c}$, there can be no such thing. This indicates that a second-order phase transition can occur here. We have depicted the scenario in 9.
(a) Symmetry breaking with the effective potential of 45. Here $T_{c}\sim 14.77{\rm K}$ and $T\sim 1{\rm K}$.
(b) Symmetry restoration with $T>T_{c}$.
Figure 9: Symmetry breaking and restoration of the thermal effective potential of 45 in flat spacetime with $m^{2}{<}0$.
Let us now add the tree level curvature term $\xi R\phi^{2}/2$ to see if there is any modification to the above phenomenon. We have
$V_{\rm{eff},\beta}=\frac{m^{2}\phi^{2}}{2}\left(1-\frac{T^{2}}{T^{2}_{c}}+\frac{\xi R}{m^{2}}\right)+\frac{\lambda\phi^{4}}{4!}$ (46)
This indicates a new critical temperature, say $T_{1c}$, given by
$T_{1c}=\left(1+\frac{\xi R}{m^{2}}\right)^{\frac{1}{2}}T_{c}$ (47)
showing that for $\xi R{>}0$, a reduction in the value of the critical temperature and an increment for the opposite sign.
At one loop, for $\beta\to 0$, we get after keeping only terms linear in the spacetime curvature,
$V_{\rm{eff,\beta}}=\frac{\lambda\phi^{4}}{4!}+\frac{m^{2}\phi^{2}}{2}\left[1-\frac{T^{2}}{T^{2}_{c}}+\frac{\xi R}{m^{2}}-\frac{3\lambda}{64\pi^{2}}-0.732\times\frac{\left(\frac{1}{6}-\xi\right)R\lambda}{16\pi^{2}m^{2}}\right]$ (48)
Now the critical temperature reads
$T_{1c}=\left[1+\frac{\xi R}{m^{2}}-\frac{3\lambda}{64\pi^{2}}-0.732\times\frac{\left(\frac{1}{6}-\xi\right)R\lambda}{16\pi^{2}m^{2}}\right]^{\frac{1}{2}}T_{c}$ (49)
For the de Sitter spacetime, $(R=4\Lambda)$, we have plotted the variation of $T_{1c}$ with respect to the quartic coupling in 10.
Figure 10: Variation of the effective critical temperature with respect to $\lambda$, 49 in the de Sitter spacetime. We have taken $\xi\sim 0.125$ and $\Lambda\sim 10^{-52}{\rm m}^{-2}$.
This completes our brief discussion on the case of negative rest mass squired of the scalar field. We wish to discuss below the usual case when $m^{2}$ is positive.
$m^{2}>0$ case :
Since 35 is essentially a high temperature expansion, and we are not setting the mass scale to be equal to $\beta^{-1}$, we will ignore the zero temperature part in 35. In particular, we have explicitly verified that even if we keep them, they bring in no qualitative changes to the symmetry restoration phenomenon at high temperature. Also, in the finite temperature part of the effective action, we have not included the odd power terms of $m_{1}$, as they originate from the zero mode fluctuations (i.e. the $n=0$ term in 28). We refer our reader to [7, 67] and references therein in this context. Nevertheless, we also have explicitly verified that for $\bar{\eta}=0$, inclusion of such terms do not alter any qualitative feature of the symmetry restoration phenomenon.
Let us accordingly begin with the thermal effective potential with zero background spacetime curvature found from 35,
$\frac{V_{\rm{eff},\beta}}{m^{4}}=\frac{\bar{\phi}^{2}}{2}+\frac{\bar{\eta}\bar{\phi}^{3}}{3!}+\frac{\lambda\bar{\phi}^{4}}{4!}+\left[\frac{\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}}{24m^{2}\beta^{2}}-0.001\times\left(2\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)+\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{2}\right)\ln{\left((m\beta)^{2}\left(1+\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)\right)}\right.$
$\left.+0.00004\times(m\beta)^{2}\left(3\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)+3\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{2}+\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{3}\right)\right.$
$\left.-4.37\times 10^{-8}(m\beta)^{4}\left(6\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{2}+4\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)+4\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{3}+\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{4}\right)\right.$
$\left.+5.38\times 10^{-10}(m\beta)^{6}\left(5\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)+10\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{2}+10\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{3}+5\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{4}+\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{5}\right)\right]$
which we have plotted in 11. Comparing this with 3 for zero temperature, we see no qualitative changes. This is expected, as if there is no symmetry breaking at zero temperature, one usually does not expect the same at finite temperature.
(a) One loop $\eta\phi^{3}+\lambda\phi^{4}$ interaction
(b) One loop $\lambda\phi^{4}$ interaction
Figure 11: Plot of the effective potential at finite temperature$(\beta=10^{-5},\eta=0.447,\lambda=0.1)$ with vanishing background curvature, 50. See main text for discussion.
For the Schwarzschild spacetime, we have from 35
$\frac{V_{\rm{eff},\beta}}{m^{4}}=\frac{\bar{\phi}^{2}}{2}+\frac{\bar{\eta}\bar{\phi}^{3}}{3!}+\frac{\lambda\bar{\phi}^{4}}{4!}+\left[\frac{1}{24}\frac{\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}}{m^{2}\beta^{2}}-0.001\times\left(2\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)+\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{2}\right)\ln\left((m\beta)^{2}\left(1+\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)\right)\right.$
$\left.+0.00004\times(m\beta)^{2}\left(3\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)+3\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{2}+\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{3}\right)\right.$
$\left.-4.37\times 10^{-8}(m\beta)^{4}\left(6\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{2}+4\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)+4\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{3}+\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{4}\right)\right.$
$\left.+5.38\times 10^{-10}(m\beta)^{6}\left(5\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)+10\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{2}+10\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{3}+5\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{4}+\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{5}\right)\right]$
$-\frac{0.0002}{m^{4}G^{4}M^{4}}\left[-0.003\times(m\beta)^{2}\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)+0.00004\times(m\beta)^{4}\left(2\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)+\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{2}\right)\right.$
$\left.+0.00001\times(m\beta)^{6}\left(3\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)+3\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{2}+\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{3}\right)\right]$ (51)
which we have plotted in 12 for $GM\sim 10^{2}{\rm m}$. Comparing them with the results of 5, we see, expectedly, no qualitatively new feature is added.
Finally, for the de Sitter spacetime, we have
$\frac{V_{\rm{eff},\beta}}{m^{4}}=\frac{\bar{\phi}^{2}}{2}+\frac{\bar{\eta}\bar{\phi}^{3}}{3!}+\frac{\lambda\bar{\phi}^{4}}{4!}+2\xi\Lambda\bar{\phi}^{2}+\left[\frac{\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}}{24m^{2}\beta^{2}}-0.001\times\left(2\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)+\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{2}\right)\ln\left((m\beta)^{2}\left(1+\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)\right)\right.$
$\left.+0.00004\times(m\beta)^{2}\left(3\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)+3\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{2}+\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{3}\right)\right.$
$\left.-4.37\times 10^{-8}(m\beta)^{4}\left(6\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{2}+4\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)+4\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{3}+\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{4}\right)\right.$
$\left.+5.38\times 10^{-10}(m\beta)^{6}\left(5\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)+10\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{2}+10\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{3}+5\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{4}+\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{5}\right)\right]$
$-0.12\left(\frac{1}{6}-\xi\right)\frac{\Lambda}{m^{2}}\left[-0.1\ln\left(m^{2}\beta^{2}\left(1+\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)\right)+0.001\times(m\beta)^{2}\left(2\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)+\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{2}\right)\right.$
$\left.+0.00001\times(m\beta)^{4}\left(3\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)+3\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{2}+\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{3}\right)\right.$
$\left.+0.00002\times(m\beta)^{6}\left(6\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{2}+4\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)+4\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{3}+\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{4}\right)\right]$
$-\frac{0.03}{m^{4}}\left(\left(\frac{1}{6}-\xi\right)^{2}16\Lambda^{2}-\frac{2}{135}\Lambda^{2}\right)\left[-0.003\times(m\beta)^{2}\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)+0.00004\times(m\beta)^{4}\left(2\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)+\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{2}\right)\right.$
$\left.+0.00001\times(m\beta)^{6}\left(3\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)+3\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{2}+\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{3}\right)\right]+0.24\times\left(\frac{1}{6}-\xi\right)^{2}\frac{\Lambda^{2}}{m^{4}}\left[-0.003\times(m\beta)^{2}\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)\right.$
$\left.+0.00004\times(m\beta)^{4}\left(2\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)+\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{2}\right)+0.00001\times(m\beta)^{6}\left(3\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)+3\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{2}+\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{3}\right)\right]$
The variation of the above effective potential can be seen in 13. We have taken $\Lambda\sim 10^{-52}{\rm m}^{-2}$ and $\xi=0.125$. Comparing this with the zero temperature result of 5, we see symmetry restoration. This must be a second order phase transition.
(a) One loop $\eta\phi^{3}+\lambda\phi^{4}$ interaction (Schwarzschild)
(b) One loop $\lambda\phi^{4}$ interaction (Schwarzschild)
Figure 12: Variation of the effective potential for one loop at finite temperature for the Schwarzschild spacetime, 51.
(a) One loop $\eta\phi^{3}+\lambda\phi^{4}$ interaction (de Sitter)
(b) One loop $\lambda\phi^{4}$ interaction (de Sitter)
Figure 13: Plot of the finite temperature effective potential in the de Sitter background ($\beta=10^{-5},\eta=0.447,\lambda=0.1,\Lambda\sim 10^{-52}{\rm m}^{-2}$ and $\xi=0.125$). Comparing this with the zero temperature result of 5, we see symmetry restoration effect at finite temperature and hence indication of a second order phase transition.
6 Two loop effective potential in curved spacetime
We now wish to extend some of our earlier results to two loop order. The effective action at this order reads ($\hbar=1$)
$-\frac{\lambda}{8}\int\sqrt{-g}d^{d}xG^{2}(x,x)+\frac{1}{12}\int\sqrt{-g}\sqrt{-g^{\prime}}d^{d}xd^{d}x^{\prime}(\eta+\lambda\phi)^{2}G^{3}(x,x^{\prime})+\ {\rm counterterm~{}contributions}$ (53)
Note that the second or the sunset integral is quadratic in the couplings. Thus we may expect a priori that the first, i.e. the double bubble integral to make the leading contributions at two loop, also supported by our earlier argument of 39. The corresponding 1PI vacuum diagrams can be seen 14.
(a) Double Bubble
(b) Sunset
(c) One loop counterterm
Figure 14: Two loop 1PI vacuum diagrams for $\lambda\phi^{4}+\eta\phi^{3}$ interactions.
At zero temperature, computation of the divergent part of the effective action and subsequent renormalisation can be seen in [4]. We also refer our reader to [40] for a renormalisation group derivation of a two loop effective potential with quartic self interaction under weak curvature approximation at zero temperature.
Let us now compute the diagrams of 14. In order to keep our expressions simple, we shall also make a linear in curvature approximation in the following. We evaluate the zero temperature effective potential first, show the renormalisation and then turn on the temperature. We will see that the latter is divergence free, as expected. Finally we shall present the finite part of the ${\cal O}(\lambda)$ double bubble diagram at finite temperature up to quadratic order in the curvature.
The double bubble:
The double bubble diagram is given by the first of 14. We have at $d=4-\epsilon$,
$-\frac{\lambda}{8}\left(\int\frac{d^{d}k}{(2\pi)^{d}}\left(\frac{1}{k^{2}+m^{2}_{1}}+\frac{\left(\frac{1}{6}-\xi\right)R}{(k^{2}+m^{2}_{1})^{2}}\right)\right)^{2}=-\frac{\lambda m^{4}_{1}}{2(4\pi)^{4}}\left(\frac{\mu^{-2\epsilon}}{\epsilon^{2}}-\frac{\mu^{-\epsilon}}{\epsilon}\left(\ln{\frac{m^{2}_{1}}{4\pi\mu^{2}}}-\psi(2)\right)-\psi(2)\ln{\frac{m^{2}_{1}}{4\pi\mu^{2}}}\right.$
$\left.+\frac{1}{2}\left(\ln{\frac{m^{2}_{1}}{4\pi\mu^{2}}}\right)^{2}+\frac{1}{4}\left(\frac{\pi^{2}}{3}+2\psi^{2}(2)-\psi^{\prime}(2)\right)\right)+\frac{\lambda\left(\frac{1}{6}-\xi\right)Rm^{2}_{1}}{4(4\pi)^{4}}\left(\frac{4\mu^{-2\epsilon}}{\epsilon^{2}}+\frac{2\mu^{-\epsilon}}{\epsilon}\left(\psi(1)+\psi(2)\right)-\frac{4\mu^{-\epsilon}}{\epsilon}\ln{\frac{m^{2}_{1}}{4\pi\mu^{2}}}\right.$
$\left.+2\left(\ln{\frac{m^{2}_{1}}{4\pi\mu^{2}}}\right)^{2}+\psi(1)\psi(2)+\frac{1}{2}\left(\frac{\pi^{2}}{3}+\psi^{2}(2)-\psi^{\prime}(2)\right)+\frac{1}{2}\left(\frac{\pi^{2}}{3}+\psi^{2}(1)-\psi^{\prime}(1)\right)+2\ln{\frac{4\pi\mu^{2}}{m^{2}_{1}}}\left(\psi(1)+\psi(2)\right)\right)$
where as we have stated earlier, we have kept terms only linear in curvature and
$\psi^{\prime}(x)=\frac{\Gamma^{\prime\prime}(x)}{\Gamma(x)}-\frac{(\Gamma^{\prime}(x))^{2}}{\Gamma^{2}(x)}$
The sunset:
The sunset integral, up to linear in $R$ terms reads
$-\frac{\left(\eta+\lambda\phi\right)^{2}m^{2}_{1}}{2(4\pi)^{4}}\left(\frac{\mu^{-2\epsilon}}{\epsilon^{2}}+\frac{\mu^{-\epsilon}}{\epsilon}\left(\frac{3}{2}+\psi(1)-\ln{\frac{m^{2}_{1}}{4\pi\mu^{2}}}\right)+2\left(\ln\frac{4\pi\mu^{2}}{m^{2}_{1}}\right)^{2}+2\left(\psi^{\prime}(1)+\psi^{2}(1)\right)\right.$
$\left.+4\psi(1)\ln\frac{4\pi\mu^{2}}{m^{2}_{1}}+\ln\frac{4\pi\mu^{2}}{m^{2}_{1}}+\psi(1)-\frac{1}{2}\right)$
$+\frac{\left(\eta+\lambda\phi\right)^{2}}{2}\frac{\left(\frac{1}{6}-\xi\right)R}{(4\pi)^{4}}\left(\frac{\mu^{-2\epsilon}}{\epsilon^{2}}+\frac{\mu^{-\epsilon}}{\epsilon}\left(\frac{1}{2}+\psi(1)-\ln{\frac{m^{2}_{1}}{4\pi\mu^{2}}}\right)+2\left(\ln\frac{4\pi\mu^{2}}{m^{2}_{1}}\right)^{2}+2\left(\psi^{\prime}(1)+\psi^{2}(1)\right)\right.$
$\left.+4\psi(1)\ln\frac{4\pi\mu^{2}}{m^{2}_{1}}+\ln\frac{4\pi\mu^{2}}{m^{2}_{1}}+\psi(1)-\frac{1}{2}\right)$ (55)
One loop counterterm contribution:
The one loop counterterm contribution reads
$\frac{1}{2}\left(\delta m^{2}+\frac{1}{2}\delta\lambda\phi^{2}+\delta\eta\phi+\delta\xi R\right)\times\int\frac{d^{d}k}{(2\pi)^{d}}\left(\frac{1}{k^{2}+m^{2}_{1}}+\frac{\left(\frac{1}{6}-\xi\right)R}{(k^{2}+m^{2}_{1})^{2}}\right)$ (56)
where $\delta m^{2}$, $\delta\lambda$ and $\delta\xi$ can be seen in 18. We next add the divergent terms coming from 54, 55, 56, in order to find out the divergence of the effective potential
$\mathbb{Div}V_{\rm{eff}}^{2-{\rm loop}}=\frac{\mu^{-2\epsilon}}{2(4\pi)^{4}\epsilon^{2}}\left(2m^{4}\lambda+m^{2}\eta^{2}+4m^{2}\eta\lambda\phi+\eta^{3}\phi+4m^{2}\lambda^{2}\phi^{2}+\frac{7}{2}\lambda\eta^{2}\phi^{2}+3\eta\lambda^{2}\phi^{3}+\frac{3}{4}\lambda^{3}\phi^{4}-2m^{2}\lambda\left(\frac{1}{6}-\xi\right)R\right.$
$\left.-\eta^{2}\left(\frac{1}{6}-\xi\right)R-4\eta\lambda\phi\left(\frac{1}{6}-\xi\right)R-\lambda^{2}\phi^{2}\left(\frac{1}{6}-\xi\right)R\right)$
$+\frac{\mu^{-\epsilon}}{2(4\pi)^{4}\epsilon}\left(\lambda\eta^{2}\phi^{2}\left(\frac{5}{2}\psi(2)-\frac{5}{2}\psi(1)-\frac{15}{4}\right)+2m^{2}\lambda^{2}\phi^{2}\left(\frac{3}{4}\psi(2)-\frac{1}{2}\psi(1)-\frac{3}{4}\right)+m^{2}\eta^{2}\left(\psi(2)-\psi(1)-\frac{3}{2}\right)\right.$
$\left.+m^{2}\eta\lambda\phi\left(2\psi(2)-2\psi(1)-3\right)+\eta^{3}\phi\left(\psi(2)-\psi(1)-\frac{3}{2}\right)+\frac{1}{2}\lambda^{3}\phi^{4}\left(\psi(2)-\psi(1)-\frac{3}{2}\right)\right.$
$\left.+\eta\lambda^{2}\phi^{3}\left(2\psi(2)-2\psi(1)-3\right)+\frac{\eta^{2}}{2}\left(\frac{1}{6}-\xi\right)R+\eta\lambda\phi\left(\frac{1}{6}-\xi\right)R+\frac{\lambda^{2}\phi^{2}}{2}\left(\frac{1}{6}-\xi\right)R\right)$ (57)
The above equation leads to the two loop counterterms
$\delta\Lambda=-\frac{\mu^{-2\epsilon}\left(m^{4}\lambda+2m^{2}\eta^{2}\right)}{2(4\pi)^{4}\epsilon^{2}}-\frac{2\mu^{-\epsilon}m^{2}\eta^{2}\left(\psi(2)-\psi(1)-\frac{3}{2}\right)}{2(4\pi)^{4}\epsilon}$
$\delta\kappa=\frac{\mu^{-2\epsilon}\left(2m^{2}\lambda+\eta^{2}\right)}{2(4\pi)^{4}\epsilon^{2}}\left(\frac{1}{6}-\xi\right)-\frac{\eta^{2}\mu^{-\epsilon}}{4(4\pi)^{4}\epsilon}\left(\frac{1}{6}-\xi\right)$
$\delta m^{2}=\frac{\left(4m^{2}\lambda^{2}+\frac{7}{2}\lambda\eta^{2}\right)\mu^{-2\epsilon}}{(4\pi)^{4}\epsilon^{2}}+\frac{2m^{2}\lambda^{2}\mu^{-\epsilon}}{(4\pi)^{4}\epsilon}\left(\frac{3}{4}\psi(2)-\frac{1}{2}\psi(1)-\frac{3}{4}\right)+\frac{\eta^{2}\lambda\mu^{-\epsilon}}{(4\pi)^{4}\epsilon}\left(\frac{5}{2}\psi(2)-\frac{5}{2}\psi(1)-\frac{15}{4}\right)$
$\delta\xi=-\frac{\mu^{-2\epsilon}\lambda^{2}}{(4\pi)^{4}\epsilon^{2}}\left(\frac{1}{6}-\xi\right)+\frac{\lambda^{2}\mu^{-\epsilon}}{2(4\pi)^{4}\epsilon}\left(\frac{1}{6}-\xi\right)$
$\delta\tau=\frac{\left(4m^{2}\eta\lambda+\eta^{3}\right)\mu^{-2\epsilon}}{2(4\pi)^{4}\epsilon^{2}}+\frac{m^{2}\eta\lambda\mu^{-\epsilon}}{2(4\pi)^{4}\epsilon}\left(2\psi(2)-2\psi(1)-3\right)+\frac{\eta^{3}\mu^{-\epsilon}}{2(4\pi)^{4}\epsilon}\left(\psi(2)-\psi(1)-\frac{3}{2}\right)$
$\delta\gamma=-\frac{4\eta\lambda\mu^{-2\epsilon}}{2(4\pi)^{4}\epsilon^{2}}\left(\frac{1}{6}-\xi\right)-\frac{\eta\lambda\mu^{-\epsilon}}{2(4\pi)^{4}\epsilon}\left(\frac{1}{6}-\xi\right)$
$\delta\eta=\frac{9\mu^{-2\epsilon}\eta\lambda^{2}}{(4\pi)^{4}\epsilon^{2}}+\frac{3\mu^{-\epsilon}\eta\lambda^{2}}{(4\pi)^{4}\epsilon}\left(2\psi(2)-2\psi(1)-3\right)$
$\delta\lambda=\frac{9\lambda^{3}\mu^{-2\epsilon}}{(4\pi)^{4}\epsilon^{2}}+\frac{6\lambda^{3}\mu^{-\epsilon}}{(4\pi)^{4}\epsilon}\left(\psi(2)-\psi(1)-\frac{3}{2}\right)$ (58)
Apart from some multiplicative numerical factors, the above result agrees with that of [4].
58 renormalises the effective potential completely, giving us its finite parts coming from the double bubble and the sunset diagrams
$V_{\rm{eff}}^{2-{\rm loop}}=-\frac{\lambda m^{4}_{1}}{2(4\pi)^{4}}\left(-\psi(2)\ln{\frac{m^{2}_{1}}{4\pi\mu^{2}}}+\frac{1}{2}\left(\ln{\frac{m^{2}_{1}}{4\pi\mu^{2}}}\right)^{2}+\frac{1}{4}\left(\frac{\pi^{2}}{3}+2\psi^{2}(2)-\psi^{\prime}(2)\right)\right)$
$+\frac{\lambda\left(\frac{1}{6}-\xi\right)Rm^{2}_{1}}{4(4\pi)^{4}}\left(2\left(\ln{\frac{m^{2}_{1}}{4\pi\mu^{2}}}\right)^{2}+\psi(1)\psi(2)+\frac{1}{2}\left(\frac{\pi^{2}}{3}+\psi^{2}(2)-\psi^{\prime}(2)\right)+\frac{1}{2}\left(\frac{\pi^{2}}{3}+\psi^{2}(1)-\psi^{\prime}(1)\right)\right.$
$\left.+2\ln{\frac{4\pi\mu^{2}}{m^{2}_{1}}}\left(\psi(1)+\psi(2)\right)\right)-\frac{\left(\eta+\lambda\phi\right)^{2}m^{2}_{1}}{2(4\pi)^{4}}\left(2\left(\ln\frac{4\pi\mu^{2}}{m^{2}_{1}}\right)^{2}+2\left(\psi^{\prime}(1)+\psi^{2}(1)\right)+4\psi(1)\ln\frac{4\pi\mu^{2}}{m^{2}_{1}}\right.$
$\left.+\ln\frac{4\pi\mu^{2}}{m^{2}_{1}}+\psi(1)-\frac{1}{2}\right)+\frac{\left(\eta+\lambda\phi\right)^{2}\left(\frac{1}{6}-\xi\right)R}{2(4\pi)^{4}}\left(2\left(\ln\frac{4\pi\mu^{2}}{m^{2}_{1}}\right)^{2}+2\left(\psi^{\prime}(1)+\psi^{2}(1)\right)+4\psi(1)\ln\frac{4\pi\mu^{2}}{m^{2}_{1}}\right.$
$\left.+\ln\frac{4\pi\mu^{2}}{m^{2}_{1}}+\psi(1)-\frac{1}{2}\right)$ (59)
As we mentioned at the beginning of this section, the ${\cal O}(\lambda)$ double bubble diagram should make the leading contribution compared to the sunset which is quadratic in the couplings. Thus if we consider only the first, we have the two loop effective potential
$V_{\rm{eff}}^{2-{\rm loop}}=V_{\rm{eff}}^{1-{\rm loop}}-\frac{\lambda m^{4}_{1}}{2(4\pi)^{4}}\left(-\psi(2)\ln{\frac{m^{2}_{1}}{4\pi\mu^{2}}}+\frac{1}{2}\left(\ln{\frac{m^{2}_{1}}{4\pi\mu^{2}}}\right)^{2}+\frac{1}{4}\left(\frac{\pi^{2}}{3}+2\psi^{2}(2)-\psi^{\prime}(2)\right)\right)$
$+\frac{\lambda\left(\frac{1}{6}-\xi\right)Rm^{2}_{1}}{4(4\pi)^{4}}\left(2\left(\ln{\frac{m^{2}_{1}}{4\pi\mu^{2}}}\right)^{2}+\psi(1)\psi(2)+\frac{1}{2}\left(\frac{\pi^{2}}{3}+\psi^{2}(2)-\psi^{\prime}(2)\right)+\frac{1}{2}\left(\frac{\pi^{2}}{3}+\psi^{2}(1)-\psi^{\prime}(1)\right)\right.$
$\left.+2\ln{\frac{4\pi\mu^{2}}{m^{2}_{1}}}\left(\psi(1)+\psi(2)\right)\right)+{\cal O}(\lambda^{2})$ (60)
where the one loop effective potential is given by 20. However, we note that the one loop effective potential contains terms quadratic in the curvature as well, whereas in the two loop computation, we have retained terms only linear in the curvature. Hence we shall add with 60 the second order in curvature contribution coming from the ${\cal O}(\lambda)$ double bubble, reading
$V_{{\rm eff},{\cal O}(R^{2})}^{2-{\rm loop}}=-\frac{\lambda}{8(4\pi)^{4}}\left({2f_{1}}\left(\ln{\frac{m^{2}_{1}}{4\pi\mu^{2}}-\psi(2)}\right)+\left(\frac{1}{6}-\xi\right)^{2}R^{2}\left(4\psi(1)\ln{\frac{4\pi\mu^{2}}{m_{1}^{2}}}+\frac{3}{2}\left(\ln{\frac{4\pi\mu^{2}}{m_{1}^{2}}}\right)^{2}+\frac{\pi^{2}}{3}+2\psi^{2}(1)-\psi^{\prime}(1)\right)\right)$ (61)
We shall not go into the detail of the derivation for the above. The above contribution only required some gravitational counterterms for the sake of renormalisation. We have plotted 60 plus 61 in 15 for the de Sitter spacetime. Note that compared to the zero temperature one loop result, 43, there is no qualitatively new change. In particular for $\bar{\eta}=0$, the symmetry breaking feature remains intact.
(a) $\eta\phi^{3}+\lambda\phi^{4}$ interaction
(b) $\lambda\phi^{4}$ interaction
Figure 15: The two loop double bubble effective potential, 60, at vanishing temperature in the de Sitter spacetime for values $\bar{\eta}=0.447,\ \lambda=0.1,\ \xi=0.125,\ \Lambda=10^{-52}{\rm m}^{-2}$. Note that for the quartic self interaction, the symmetry breaking feature remains intact compared to that of one loop, 5.
6.1 Two loop at finite temperature
Finally, let us come to finite temperature computations at two loop. In order to demonstrate the feature of renormalisability clearly, instead of going to discretisation of the Euclidian $k^{0}$ as of 4, we make the following decomposition (e.g. [23, 26] and references therein)
$\frac{1}{k^{2}+m^{2}}\rightarrow\frac{1}{k^{2}+m^{2}}+2\pi n_{B}({|k^{0}|})\delta(k^{2}-m^{2})=\frac{1}{k^{2}+m^{2}}+n_{1}(\beta,k)\qquad({\rm say})$ (62)
And,
$\frac{1}{\left(k^{2}+m^{2}\right)^{2}}\rightarrow\frac{1}{\left(k^{2}+m^{2}\right)^{2}}+n_{2}(\beta,k)$ (63)
where $n_{B}({|k^{0}|})=(\exp{\beta|k^{0}|}-1)^{-1}$ is the Bose distribution function.
Thus the integrations will be like,
$\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{k^{2}+m^{2}}\rightarrow\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{k^{2}+m^{2}}+S_{1}(\beta)=-m^{2}\left(\frac{2\mu^{-\epsilon}}{\epsilon}+\psi(2)+\ln{\frac{4\pi\mu^{2}}{m^{2}}}\right)+S_{1}(\beta)$ (64)
where $S_{1}(\beta)$ is temperature dependent and finite. Likewise
$\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{(k^{2}+m^{2})^{2}}\rightarrow\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{(k^{2}+m^{2})^{2}}+S_{2}(\beta)=\left(\frac{2\mu^{-\epsilon}}{\epsilon}+\psi(1)+\ln{\frac{4\pi\mu^{2}}{m^{2}}}\right)+S_{2}(\beta)$ (65)
Double bubble at finite temperature:
The double bubble integral at finite temperature and up to ${\cal O}(R)$ reads
$-\frac{\lambda}{8}\left(\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{k^{2}+m^{2}_{1}}+S_{1}(\beta)+\left(\frac{1}{6}-\xi\right)R\left(\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{(k^{2}+m^{2}_{1})^{2}}+S_{2}(\beta)\right)\right)^{2}$
$=-\frac{\lambda}{8}\left(\left(\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{k^{2}+m^{2}_{1}}\right)^{2}+2\left(\frac{1}{6}-\xi\right)R\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{k^{2}+m^{2}_{1}}\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{(k^{2}+m^{2}_{1})^{2}}\right)$
$-\frac{\lambda}{8}\left(-\left(m^{2}+\eta\phi+\frac{1}{2}\lambda\phi^{2}\right)\frac{4\mu^{-\epsilon}}{(4\pi)^{2}\epsilon}S_{1}(\beta)-\left(\frac{1}{6}-\xi\right)R\left(m^{2}+\eta\phi+\frac{1}{2}\lambda\phi^{2}\right)\frac{4\mu^{-\epsilon}}{(4\pi)^{2}\epsilon}S_{2}(\beta)+\left(\frac{1}{6}-\xi\right)R\frac{4\mu^{-\epsilon}}{(4\pi)^{2}\epsilon}S_{1}(\beta)\right)$
$+\rm{Finite\,\,terms}$ (66)
Note that the first two terms are temperature independent divergent terms whereas the next three terms are temperature dependent divergent terms.
Sunset at finite temperature:
The feynmann integral for Sunset diagram will be
$\frac{\left(\eta+\lambda\phi\right)^{2}}{12}\left(\int\frac{d^{d}k_{1}}{(2\pi)^{d}}\frac{d^{d}k_{2}}{(2\pi)^{d}}\left(\frac{1}{k_{1}^{2}+m^{2}_{1}}+n_{1}(k_{1})\right)\left(\frac{1}{k_{2}^{2}+m^{2}_{1}}+n_{1}(\beta,k_{2})\right)\left(\frac{1}{(k_{1}+k_{2})^{2}+m^{2}_{1}}+n_{1}(\beta,k_{1}+k_{2})\right)\right.$
$\left.+\left(\frac{1}{6}-\xi\right)R\int\frac{d^{d}k_{1}}{(2\pi)^{d}}\frac{d^{d}k_{2}}{(2\pi)^{d}}\left(\frac{1}{k_{1}^{2}+m^{2}_{1}}+n_{1}(\beta,k_{1})\right)\left(\frac{1}{(k_{2}^{2}+m^{2}_{1})^{2}}+n_{2}(k_{2})\right)\left(\frac{1}{(k_{1}+k_{2})^{2}+m^{2}_{1}}+n_{1}(\beta,k_{1}+k_{2})\right)\right.$
$\left.+\left(\frac{1}{6}-\xi\right)R\int\frac{d^{d}k_{1}}{(2\pi)^{d}}\frac{d^{d}k_{2}}{(2\pi)^{d}}\left(\frac{1}{\left(k_{1}^{2}+m^{2}_{1}\right)^{2}}+n_{2}(\beta,k_{1})\right)\left(\frac{1}{k_{2}^{2}+m^{2}_{1}}+n_{1}(\beta,k_{2})\right)\left(\frac{1}{(k_{1}+k_{2})^{2}+m^{2}_{1}}+n_{1}(\beta,k_{1}+k_{2})\right)\right.$
$\left.+\left(\frac{1}{6}-\xi\right)R\int\frac{d^{d}k_{1}}{(2\pi)^{d}}\frac{d^{d}k_{2}}{(2\pi)^{d}}\left(\frac{1}{k_{1}^{2}+m^{2}_{1}}+n_{1}(\beta,k_{1})\right)\left(\frac{1}{k_{2}^{2}+m^{2}_{1}}+n_{1}(\beta,k_{2})\right)\left(\frac{1}{\left((k_{1}+k_{2})^{2}+m^{2}_{1}\right)^{2}}+n_{2}(\beta,k_{1}+k_{2})\right)\right)$
$=\frac{\left(\eta+\lambda\phi\right)^{2}}{12}\left(\int\frac{d^{d}k_{1}}{(2\pi)^{d}}\frac{d^{d}k_{2}}{(2\pi)^{d}}\frac{1}{k_{1}^{2}+m^{2}_{1}}\frac{1}{k_{2}^{2}+m^{2}_{1}}\frac{1}{(k_{1}+k_{2})^{2}+m^{2}_{1}}\right.$
$\left.+\left(\frac{1}{6}-\xi\right)R\int\frac{d^{d}k_{1}}{(2\pi)^{d}}\frac{d^{d}k_{2}}{(2\pi)^{d}}\frac{1}{k_{1}^{2}+m^{2}_{1}}\frac{1}{(k_{2}^{2}+m^{2}_{1})^{2}}\frac{1}{(k_{1}+k_{2})^{2}+m^{2}_{1}}\right.$
$\left.+\left(\frac{1}{6}-\xi\right)R\int\frac{d^{d}k_{1}}{(2\pi)^{d}}\frac{d^{d}k_{2}}{(2\pi)^{d}}\frac{1}{(k_{1}^{2}+m^{2}_{1})^{2}}\frac{1}{k_{2}^{2}+m^{2}_{1}}\frac{1}{(k_{1}+k_{2})^{2}+m^{2}_{1}}\right.$
$\left.+\left(\frac{1}{6}-\xi\right)R\int\frac{d^{d}k_{1}}{(2\pi)^{d}}\frac{d^{d}k_{2}}{(2\pi)^{d}}\frac{1}{k_{1}^{2}+m^{2}_{1}}\frac{1}{k_{2}^{2}+m^{2}_{1}}\frac{1}{((k_{1}+k_{2})^{2}+m^{2}_{1})^{2}}\right)$
$+\frac{\left(\eta+\lambda\phi\right)^{2}}{12}\left(\frac{6\mu^{-\epsilon}}{\epsilon}S_{1}(\beta)+\left(\frac{1}{6}-\xi\right)R\frac{6\mu^{-\epsilon}}{\epsilon}S_{2}(\beta)\right)+\rm{Finite\,\,terms}$ (67)
The first four terms in the final expression are temperature independent divergent terms whereas next two terms are temperature dependent divergent terms.
One loop counterterm contribution at finite temperature:
The relevant integral reads
$\frac{1}{2}\left(\delta m^{2}+\frac{1}{2}\delta\lambda\phi^{2}+\delta\eta\phi+\delta\xi R\right)\left(\int\frac{d^{d}k}{(2\pi)^{d}}\left(\frac{1}{k^{2}+m^{2}_{1}}+\frac{\left(\frac{1}{6}-\xi\right)R}{(k^{2}+m^{2}_{1})^{2}}\right)+S_{1}(\beta)+\left(\frac{1}{6}-\xi\right)RS_{2}(\beta)\right)$
$=\frac{1}{2}\left(\delta m^{2}+\frac{1}{2}\delta\lambda\phi^{2}+\delta\eta\phi+\delta\xi R\right)\left(\int\frac{d^{d}k}{(2\pi)^{d}}\left(\frac{1}{k^{2}+m^{2}_{1}}+\frac{\left(\frac{1}{6}-\xi\right)R}{(k^{2}+m^{2}_{1})^{2}}\right)\right)$
$+\frac{\mu^{-\epsilon}}{2(4\pi)^{2}\epsilon}\left(-m^{2}\lambda-\eta^{2}-\frac{3}{2}\lambda^{2}\phi^{2}-3\eta\lambda+\left(\frac{1}{6}-\xi\right)\lambda R\right)\left(S_{1}(\beta)+\left(\frac{1}{6}-\xi\right)RS_{2}(\beta)\right)$ (68)
We now collect and add all the divergent terms from 66, 67 and 68. We obtain after some calculations,
${\rm Div}V_{\rm{eff},\beta}^{2-{\rm loop}}=-\frac{\lambda}{8}\left(\left(\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{k^{2}+m^{2}_{1}}\right)^{2}+2\left(\frac{1}{6}-\xi\right)R\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{k^{2}+m^{2}_{1}}\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{(k^{2}+m^{2}_{1})^{2}}\right)$
$+\frac{\left(\eta+\lambda\phi\right)^{2}}{12}\left(\int\frac{d^{d}k_{1}}{(2\pi)^{d}}\frac{d^{d}k_{2}}{(2\pi)^{d}}\frac{1}{k_{1}^{2}+m^{2}_{1}}\frac{1}{k_{2}^{2}+m^{2}_{1}}\frac{1}{(k_{1}+k_{2})^{2}+m^{2}_{1}}\right.$
$\left.+\left(\frac{1}{6}-\xi\right)R\int\frac{d^{d}k_{1}}{(2\pi)^{d}}\frac{d^{d}k_{2}}{(2\pi)^{d}}\frac{1}{k_{1}^{2}+m^{2}_{1}}\frac{1}{(k_{2}^{2}+m^{2}_{1})^{2}}\frac{1}{(k_{1}+k_{2})^{2}+m^{2}_{1}}\right.$
$\left.+\left(\frac{1}{6}-\xi\right)R\int\frac{d^{d}k_{1}}{(2\pi)^{d}}\frac{d^{d}k_{2}}{(2\pi)^{d}}\frac{1}{(k_{1}^{2}+m^{2}_{1})^{2}}\frac{1}{k_{2}^{2}+m^{2}_{1}}\frac{1}{(k_{1}+k_{2})^{2}+m^{2}_{1}}\right.$
$\left.+\left(\frac{1}{6}-\xi\right)R\int\frac{d^{d}k_{1}}{(2\pi)^{d}}\frac{d^{d}k_{2}}{(2\pi)^{d}}\frac{1}{k_{1}^{2}+m^{2}_{1}}\frac{1}{k_{2}^{2}+m^{2}_{1}}\frac{1}{((k_{1}+k_{2})^{2}+m^{2}_{1})^{2}}\right)$ (69)
which is temperature independent, showing renormalisation is independent of the temperature. This serves as a consistency check of our calculations.
After renormalisation, we can compute the finite part of the two loop effective potential at finite temperature. We wish to quote below only the result of the double bubble diagram we have found,
$V^{(2)}_{\rm{db},\beta}=V^{(2)}_{\rm{db}}+\frac{1}{576\beta^{4}}+\frac{m_{1}^{2}}{64\pi^{2}\beta^{2}}+\frac{m^{8}_{1}\beta^{4}}{65536\pi^{8}}\zeta^{2}(3)-\frac{m_{1}}{192\pi\beta^{3}}+\frac{m^{4}_{1}}{256\pi^{4}}\left(\ln\frac{m_{1}\beta}{2\pi}-\psi(1)\right)^{2}-\frac{m^{2}_{1}}{16\pi^{2}}\left(\ln\frac{m_{1}\beta}{2\pi}-\psi(1)\right)\left(\frac{1}{24\beta^{2}}\right.$
$\left.-\frac{m_{1}}{8\pi\beta}+\frac{m_{1}^{4}\beta^{2}}{256\pi^{4}}\zeta(3)-\frac{m_{1}^{6}\beta^{4}}{(32)(64)\pi^{6}}\zeta(5)+\frac{15m_{1}^{8}\beta^{6}}{(96)(256)\pi^{8}}\zeta(7)\right)+\frac{m_{1}^{4}\zeta(3)}{6144\pi^{4}}-\frac{m_{1}^{5}\beta\zeta(3)}{2048\pi^{5}}-\frac{m_{1}^{6}\beta^{2}\zeta(5)}{49152\pi^{6}}+\frac{15m_{1}^{8}\beta^{4}\zeta(7)}{589824\pi^{8}}$
$+\frac{m_{1}^{7}\beta^{3}\zeta(5)}{16384\pi^{7}}-\frac{15m_{1}^{7}\beta^{3}\zeta(7)}{196608\pi^{9}}-\frac{m_{1}^{10}\beta^{6}\zeta(3)\zeta(5)}{524288\pi^{10}}+2\left(\frac{1}{6}-\xi\right)R\left(\frac{1}{384\pi\beta m_{1}}-\frac{1}{128\pi^{2}\beta^{2}}\right.$
$\left.-\frac{m^{2}_{1}}{256\pi^{4}}\left(\ln\frac{m_{1}\beta}{2\pi}-\psi(1)\right)^{2}+\frac{m^{2}_{1}}{16\pi^{2}}\left(\ln\frac{m_{1}\beta}{2\pi}-\psi(1)\right)\left(\frac{1}{16\pi\beta m_{1}}-\frac{m_{1}^{2}\beta^{2}}{128\pi^{4}}\zeta(3)+\frac{3m_{1}^{4}\beta^{4}}{2048\pi^{6}}\zeta(5)-\frac{15m_{1}^{6}\beta^{6}}{3072\pi^{8}}\zeta(7)\right)\right.$
$\left.+\frac{1}{16\pi^{2}}\left(\ln\frac{m_{1}\beta}{2\pi}-\psi(1)\right)\left(\frac{1}{24\beta^{2}}-\frac{m_{1}}{8\pi\beta}+\frac{m_{1}^{4}\beta^{2}}{256\pi^{4}}\zeta(3)-\frac{m_{1}^{6}\beta^{4}}{(32)(64)\pi^{6}}\zeta(5)+\frac{15m_{1}^{8}\beta^{6}}{(96)(256)\pi^{8}}\zeta(7)\right)-\frac{m_{1}^{2}\zeta(3)}{3072\pi^{4}}\right.$
$\left.+\frac{m_{1}^{3}\beta\zeta(3)}{4096\pi^{5}}+\frac{m_{1}^{3}\beta\zeta(3)}{1024\pi^{5}}+\frac{3m_{1}^{4}\beta^{2}\zeta(5)}{49152\pi^{6}}-\frac{m_{1}^{5}\beta^{3}\zeta(5)}{32768\pi^{7}}-\frac{3m_{1}^{5}\beta^{3}\zeta(5)}{16384\pi^{7}}-\frac{m_{1}^{6}\beta^{4}\zeta(3)^{2}}{32768\pi^{8}}\right.$
$\left.-\frac{15m_{1}^{6}\beta^{4}\zeta(7)}{3072\pi^{8}}+\frac{15m_{1}^{7}\beta^{5}\zeta(7)}{24576\pi^{9}}-\frac{15m_{1}^{7}\beta^{5}\zeta(7)}{393216\pi^{9}}+\frac{3m_{1}^{8}\beta^{6}\zeta(3)\zeta(5)}{524288\pi^{10}}+\frac{m_{1}^{8}\beta^{6}\zeta(3)\zeta(5)}{262144\pi^{10}}\right)+{\cal O}(R^{2})$ (70)
where $V^{(2)}_{\rm{db}}$ is the corresponding zero temperature part. Inclusion of ${\cal O}(R^{2})$ terms to the above result is easy, but we shall not go into such computation here.
7 Conclusion
Let us now summarise our work. In this paper we have considered the problem of the effective potential at zero and finite temperatures in a general curved spacetime. We have retained terms up to quadratic order in the curvature in the expansion of the Feynman propagator. We have focused upon the spontaneous symmetry breaking, its restoration and phase transition phenomenon. As we have emphasised earlier as well, one of our objectives was to see whether the quadratic in curvature terms can bring in some qualitatively new physical effect.
We have computed in 3, the one loop effective potential at zero temperature, up to the quadratic order of the spacetime curvature for quartic and cubic self interactions. In 4, we extended these results at finite temperature, with high temperature approximation. We discuss spontaneous symmetry breaking, its restoration and phase transition at zero and finite temperatures in 5, 5.1. We next generalise some of these results to two loop in 6. In particular, the contribution from the ${\cal O}(\lambda)$ two loop double bubble diagram under local approximation to the effective potential, at finite temperature has been explicitly presented, 6.1. For the de Sitter spacetime in particular, we have shown that we can have symmetry breaking for a scalar even with a positive rest mass squared and a positive non-minimal coupling, at zero temperature, at both one and two loop level. This cannot be achieved by linear curvature approximation and the result remains valid for a very large range of renormalisation scale, even though $\Lambda$ is taken to be as tiny as its current observed value.
The above results perhaps indicate that we should employ some non-perturbative techniques in order to investigate further the symmetry breaking and phase transition phenomena in curved spacetimes. Note that this not only means resummation of the coupling constant, but also at least some of the curvature terms appearing in the expansion of the propagator. This proposition is also motivated by the one loop self energy computation, 27. Similar analyses would be highly interesting for (non-)Abelian gauge theories. We hope to come back to this issue sometimes soon in a different publication.
Acknowledgement
The work of VN is supported by the junior research fellowship of University Grants Comission, Govt. of India (Ref. no. 211610068936). He also wishes to thank B. L. Hu and I. L. Shapiro for useful communications.